What is the minimum mass of water that would be needed to dissolves 45 g of Z at 62 degree celcius.



1
Expert's answer
2021-10-07T02:41:01-0400

Solubility ="\frac{mass \space of \space compound}{mass \space of solvent(water)}\times 100g"

So, when given solubility, minimum mass of water required in this case "=\frac{45}{solubility}\times 100g"